OpenStudy (anonymous):

Write the slope-intercept equation of the line that passes through the point (-6,-2) and is parallel to the line 2x + 3y = 6.

OpenStudy (anonymous):

2x+3y=6 y=? => y=-2/3*x+2 -------------------- y-(-2)=-2/3*(x-(-6)) y=-2/3*x-6

OpenStudy (anonymous):

parallel: y=-2/3x+c

OpenStudy (anonymous):

but the answer is y=-2/3x-6 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay thank you.